### v2

@@ Abstract
In this work, we focus on data visualizations because they are the most prevalent form of data representation in practice. Data visualizations that are produced by a designer for an audience tend to be designed in a way that is relatively unchangeable. As a material, we use the metaphor that the creator of a visualization manipulated their design while it was in a softer state, like clay. And eventually, the clay is hardened into a state that is presented to the user. Often visualization design artifacts cannot easily be altered by an end-user after they are created. Pixels cannot be moved, graphics cannot be re-embedded. The clay has been fired and the visualization is now baked. We intend to make the visual and non-visual characteristics of accessible data representations easier to change and manipulate for end-users. We want a material that is softer than software. But we also don’t want a fully malleable interface, like in end-user programming, either. We propose to call this space of design “softerware.”
We wanted to explore material softness with constraints. We conjecture that advancements in malleable interfaces and end-user programming are too system-centric and open-ended. End-user programming puts too much burden on end-users to know the language and symbols of interface-building in order to build their own interfaces. Instead, we chose to enable end-users to have agency over a data visualization interface by exposing a preferences-driven menu of options built from our existing knowledge of visualization and accessibility.
Exploring the softerware gap in data visualization became our primary research focus, which led us to formulate the following qualitative exploratory research questions:
@@
We contribute our findings from this research to the larger accessibility and visualization communities in hopes that we can inform and inspire future work that investigates softerware systems, end-user design, preferences-based user experiences, and fluid and malleable interfaces focused on end-users with disabilities. Our future work will focus on completing a finished version of our prototype and deploying it at scale within Highsoft’s Highcharts ecosystem.
@@
In general, accessibility concerns itself with a broad spectrum of barriers that people with different disabilities face. And while most literature focuses on visual disabilities [2], there are growing resources on areas such as cognitive/intellectual disabilities [3], neurodivergence [4], and both research and systems exploring epilepsy and vestibular/motion inaccessibility in visualization [5].
Yet despite these resources, making data visualizations more accessible remains a difficult task for practitioners [6]. Some accessibility guidelines even conflict, for example on the topic of patterns and textures used in charts. One side stresses that patterns are harmful to cognitive and visual accessibility while another stresses that redundant encoding strategies are necessary [7]. Understanding how to make the correct design decisions may sometimes be impossible. Either existing guidelines are incorrect or it is possible that access friction becomes inevitable the more we know what different barriers look like for different people with disabilities.
@@
While a venn diagram of softerware, end-user programming, and preferences-based parameter controls would have significant overlaps, softerware has distinct differences. Softerware differs from end-user programming by prioritizing accessibility-centered constraints and language that align with users’ lived experiences, rather than exposing general-purpose interface-building constructs focused on system- or task-centric language. It can be seen, largely, as a subset of some flavors of end-user programming. Additionally, softerware can be differentiated from common preferences-based parameter controls because it does not simply expose a set of configurable options, but instead frames these options through anticipated access barriers, prioritizing their discoverability, relevance, and impact on usability. Also, notably, softerware is concerned with the design of systems, where parameter-based solutions are just one possible design approach. We imagine that softerware also includes broader system design directions.
We contribute the concept of softerware to the larger community of researchers and practitioners because we argue it is a useful construct that can help us categorize past work, improve existing projects, and inspire new directions. Softerware systems have been part of existing work for decades, but we lack a cohesive way to refer to designing and engineering experiences that enable end-users to have agency over malleable interfaces without entering into the territory of end-user programming or interfaces with controllable parameters.
@@ RESULTS
The second theme from our practitioners (3) was the concern that end-users would be able to create a misleading or harmful data visualization. For example, encoding area size and aspect ratios can both be misleading or deceptive, yet being able to manipulate these for accessibility and contextual barriers (such as viewing a chart designed for desktop on a mobile phone) are important design considerations. Users may accidentally adjust features of a visualization while self-fitting that actually create problematic designs. Softerware systems must be designed with reasoned constraints to prevent harmful end-user design.
@@
The final theme from our practitioners was around authoring and design-tuning via softerware, where 3 participants discussed using a direct-manipulation or LLM-based softerware interface to author data visualizations and 1 of the 3 also mentioned that large-scale, privacy-preserving data collection from users could be used to create smarter design defaults in the future. We believe that both suggestions mirror existing work that speaks of “design-through-use” and “continuous-co-design,” demonstrating more value to users over the lifetime of a system.
@@
Our low vision users (P4, P7, P8) requested a feature to directly show what different options would look like in the preferences menu itself. For example, the “text size” options would either have a preview of the text size shown for each option in the menu or nearby (like showing “Large” in the actual resulting size). Low vision users in particular often use high levels of magnification and zoom, so the live results shown in the visualization space required users to go back and forth between the menu once an option was chosen and back into the chart space. This reduces overall effort, providing more value for faster interactions.
@@
The next set of themes that emerged were considerations that both our end-users and our visualization practitioners shared, which we frame as system-class considerations [20]. This theme emerged thanks to the accessibility knowledge and expertise of our end-users and engineering concerns of our practitioners and we believe will be especially helpful for future system and tool designers.
@@
We followed up by asking whether certain contexts would make persistence more or less important. We asked users whether a random website or news article with a chart in it would be worth their time, to which most users replied, “no.” However, P5 noted that “This is so fun that if it was there I still might play around with it and use it, especially if I had the time.” P4 related this issue to an existing frustration with video games, noting that having to set up repeated options for every game was time consuming. It would be nice if they could “do this once and forget it.”
@@
In addition, all 4 of our practitioners suggested that there would need to be a system in place, either at the operating system level or as some kind of service hosted by a platform, where these settings could be recognized and ingested. For cross-system interoperability to be made possible, it would require establishing standards for customization, standards for preserving user privacy, and coordination with the larger community of visualization toolmakers.
@@
Our first major user-centered finding was that no participant chose the same set of preferences as another. Every user discussed different reasons for justifying their choice of options. Users even chose options that others specifically emphasized were inaccessible to them. An example of this was a tension in preference for and against use of “dark mode” designs.
@@
P4 mentioned that they had “night blindness” (nyctalopia), which is why dark mode designs are helpful for them. However, P7 also mentioned that they had progressive nyctalopia, but dark mode makes an interface “virtually impossible” to them.
@@
Our blind participants also had different justifications and preferences for their text and audio customizations. Some justified their differing preferences with similar justifications, such as cognitive accessibility and text description length. For example, P9 stressed that “I prefer to keep things simple” to “avoid overwhelm” while P2 said, “more information is better than less, when it comes to data.” Both P9 and P2 preferred “accessible defaults,” but disagreed on what length of textual descriptions should be default.
@@
P6 was interested in different ways that this process could become easier, suggesting
@@
In light of our findings, visualization tool designers are encouraged to reframe their approach to accessible design by incorporating the softerware principles discussed in this work. Tool designers should consider moving beyond static, one-size-fits-all solutions and instead build platforms that support dynamic, user-driven customization. This entails incorporating accessible defaults that serve as a strong foundation while also providing intuitive, context-sensitive controls for personalization. By doing so, designers can help mitigate the inherent access friction that arises when a design optimized for one group inadvertently excludes another. Moreover, the integration of features such as persistent user profiles and live previews not only streamlines the customization process but also helps maintain consistency across devices and applications. Ultimately, embracing these strategies can bridge the gap between the accessibility and visualization communities, fostering an ecosystem where the diverse needs of users are met without compromising the integrity and clarity of data representation.
Our vision of data visualization softerware demands more involvement from research and industry. In order to offer as much value as possible to end-users, we need standards set for accessibility profiles and we need data visualization software, libraries, and applications to respect and be able to contribute to those profiles. We want to encourage researchers to investigate further the needs of people with disabilities, designers to imagine new interfaces and interaction paradigms for end users, and engineers to build robust systems that are capable of not only respecting a user’s preferences and customizations, but providing persistence, interoperability, and system-class infrastructure.

### v3

@@ Abstract
In this work, we focus on data visualizations because they are the most prevalent form of data representation in practice. Data visualizations that are produced by a designer for an audience tend to be designed in a way that is relatively unchangeable. As a material, we use the metaphor that the creator of a visualization manipulated their design while it was in a softer state, like clay. And eventually, the clay is hardened into a state that is presented to the user. Often visualization design artifacts cannot easily be altered by an end-user after they are created. Pixels cannot be moved, graphics cannot be re-embedded. The clay has been fired and the visualization is now baked. We intend to make the visual and non-visual characteristics of accessible data representations easier to change and manipulate for end-users. We want a material that is softer than software. And while malleable interfaces and end-user programming are a promising starting point for our work, their design principles are often too fluid and unconstrained. We propose to call the space of design between too-hard and too-fluid software system design “softerware.”
Exploring the softerware gap in visualization became our primary focus, which led us to formulate the following qualitative exploratory research questions:
@@
We contribute our findings from this research to the larger accessibility and visualization communities in hopes that we can inform and inspire future work that investigates softerware systems, end-user design, preferences-based user experiences, and malleable interfaces focused on end-users with disabilities. Our future work is dedicated to improving upon and deploying our prototype at scale within Highsoft’s Highcharts.
@@
In general, accessibility concerns itself with a broad spectrum of barriers that people with different disabilities face. And while most literature focuses on visual disabilities [2], there are growing resources on areas such as cognitive/intellectual disabilities [3], neurodivergence [4], and research exploring epilepsy and vestibular/motion inaccessibility in visualization [5].
Yet despite these resources, making data visualizations more accessible remains a difficult task for practitioners [6]. Some accessibility guidelines even conflict, for example on the topic of patterns and textures used in charts. One side stresses that patterns are harmful to cognitive and visual accessibility while another stresses that redundant encoding strategies are necessary [7]. Understanding how to make the correct design decisions may sometimes be impossible. Either existing guidelines are incorrect or it is possible that access friction becomes inevitable the more we know what barriers exist for different people.
@@
While a venn diagram of softerware, end-user programming, and preferences-based parameter controls would have significant overlaps, softerware has distinct differences. Both end-user programming and preferences-based parameter controls inform our work.
However, softerware differs from end-user programming by prioritizing accessibility-centered constraints and language that align with users’ lived experiences, rather than exposing general-purpose interface-building constructs focused on system- or task-centric language. It can be seen, largely, as a subset of some flavors of end-user programming.
Additionally, softerware can be differentiated from common preferences-based parameter controls because it does not simply expose a set of configurable options, but instead frames these options through anticipated access barriers, prioritizing their discoverability, relevance, and impact on usability. Also, notably, softerware is concerned with the design of systems, where parameter-based solutions are just one possible design approach. We imagine that softerware also includes broader system design directions than preferences-based parameter controls.
We contribute the concept of softerware to the larger community of researchers and practitioners because we argue it is a useful construct that can help us categorize past work, improve existing projects, and inspire new directions. Softerware systems have been part of existing work for decades, but we lack a cohesive way to refer to designing and engineering experiences that enable end-users to have agency over interfaces, especially in the field of data visualization.
@@ RESULTS
The second theme from our practitioners (3) was the concern that end-users would be able to create a misleading or harmful data visualization. For example, encoding area size and aspect ratios can both be misleading or deceptive, yet being able to manipulate these for accessibility and contextual barriers (such as viewing a chart designed for desktop on a mobile phone) are important design considerations. Softerware systems must be designed with reasoned constraints to prevent harmful end-user design.
@@
The final theme from our practitioners was around authoring and design-tuning via softerware, where 3 participants discussed using a direct-manipulation or LLM-based softerware interface to author data visualizations and 1 of the 3 also mentioned that large-scale, privacy-preserving data collection from users could be used to create smarter design defaults in the future. We believe that both suggestions would demonstrate more value to users over the lifetime of a system.
@@
Our low vision users (P4, P7, P8) requested a feature to directly show what different options would look like in the preferences menu. For example, the “text size” options would display the text size shown for each option in the menu. Low vision users in particular often use high levels of magnification and zoom, so the live results shown in the visualization space required users to go back and forth between the menu once an option was chosen and back into the chart space. In-menu previewing would reduce their overall effort, providing greater value.
@@
The next set of themes that emerged were considerations that both our end-users and our visualization practitioners shared, which we frame as system-class considerations [20]. This theme emerged thanks to the accessibility knowledge and expertise of our end-users and engineering concerns of our practitioners and we believe will be especially helpful for future system work.
@@
We followed up by asking whether certain contexts would make persistence more or less important. We asked users whether a random website or news article with a chart in it would be worth their time, to which most users replied, “no.” However, P5 noted that “This is so fun that if it was there I still might play around with it and use it, especially if I had the time.” P4 related this issue to an existing frustration with video games, noting that having to set up options for every game was time consuming. It would be nice if they could “do this once and forget it.”
@@
In addition, all 4 of our practitioners suggested that there would need to be a system in place, either at the operating system level or as some kind of service hosted by a platform, where these settings could be recognized and ingested. For cross-system interoperability to be made possible, it would require establishing standards for customization and preserving user privacy.
@@
Our first major user-centered finding was that no participant chose the same set of preferences as another. Every user discussed different reasons for justifying their choice of options. Users even chose options that others specifically emphasized were inaccessible to them. An example of this was a tension in preference for and against use of “dark mode” designs. P4 mentioned that they had “night blindness” (nyctalopia), which is why dark mode designs are helpful for them.
@@
However, P7 also mentioned that they had progressive nyctalopia, but dark mode makes an interface “virtually impossible” to them.
@@
Some of our blind participants justified their differing preferences on audio and text accessibility with similar reasoning, for example arguing for cognitive accessibility when deciding on text description length. For example, P9 stressed that “I prefer to keep things simple” to “avoid overwhelm” while P2 said, “more information is better than less, when it comes to data.” Both P9 and P2 preferred “accessible defaults,” but disagreed on what length should be default.
@@
P6 was interested in different ways that this process could become easier,
@@
LIMITATIONS
Our work has a limited scope, overall. We intentionally chose evaluate our softerware prototype with users who are blind and those who are low vision, in order to elicit access friction and justify the use of our softerware solution. However, significantly more work must be done to understand how to design softerware systems that support people with other disabilities.
Additionally, as participants in our study suggested, there are other interface design choices that are possible if there is an underlying softerware infrastructure available. Language-driven, direct manipulation, and multi-modal interfaces could all be used to explore softerware, and our work does not address how to make these interfaces possible or successful.
Lastly, our work focused on a softerware prototype for Highcharts. In order for other visualization tools to be built with softerware infrastructure, more work will need to explore scalable standards for visualization personalization.
@@
In light of our findings, visualization tool designers are encouraged to reframe their approach to accessible design by incorporating the softerware principles discussed in this work. Tool designers should consider moving beyond static, one-size-fits-all solutions and instead build platforms that support dynamic, user-driven customization. This entails incorporating accessible defaults that serve as a strong foundation while also providing intuitive, context-sensitive controls for personalization. By doing so, designers can help mitigate the inherent access friction that arises when a design optimized for one group excludes another.
Our vision of data visualization softerware demands more involvement from research and industry. We want to encourage researchers to investigate further the needs of people with disabilities, designers to imagine new softerware interfaces and interaction paradigms, and engineers to build robust systems that are capable of not only respecting a user’s preferences and customizations, but providing persistence, interoperability, and system-class infrastructure.